OpenStudy (anonymous):

A softball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 5. h = height t = time in seconds Find the maximum height of the ball. a) 13.9 feet b) 7.6 feet c) 12.2 feet d) 8.3 feet

OpenStudy (anonymous):

21 feet

OpenStudy (anonymous):

21 isnt an option.

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

how if its nt an option .??

OpenStudy (anonymous):

that i dont konw

OpenStudy (anonymous):

would any of the other answers maybe be it .?

OpenStudy (anonymous):

I got 21 too! It attains maximum when t=1

OpenStudy (anonymous):

see i told u

OpenStudy (anonymous):

That's what I got too

OpenStudy (anonymous):

well what am i supposed to do if all these are wrong .?!!?!?

OpenStudy (anonymous):

Check if you've typed out the equation correctly?

OpenStudy (anonymous):

A softball is thrown up into the air. The ball's height in feet is modeled by the equation h(t) = -16t2 + 32t + 5. h = height t = time in seconds Find the maximum height of the ball. a) 13.9 feet b) 7.6 feet c) 12.2 feet d) 8.3 feet thats the question

OpenStudy (anonymous):

options are wrong or the question is

OpenStudy (anonymous):

with ur question the answer is 21

OpenStudy (anonymous):

aw man :/

OpenStudy (anonymous):

well since all this is wrong can one of you help me with another problem

OpenStudy (anonymous):

There's no "none of the above" or "other?"

OpenStudy (anonymous):

nope im jt gna do another problem do you mind helping me.?

myininaya (myininaya):

h'(t)=-32t+32 -32t+32=0 t=1 h(1)=-16+32+5=16+5=21 gj him

OpenStudy (anonymous):

this is crazy. . .